OpenStudy (anonymous):

The axis of symmentry for y2- 6y + 9 -x = 0

OpenStudy (anonymous):

yes i got that too but sadly it isn't an answer choice the answer choices are y=1 y=2 y=3 y=4

OpenStudy (anonymous):

y2- 6y + 9 -x = 0 (y-3)^2= x (y-3) =0 (y-3)=0 y=3

OpenStudy (anonymous):

thank you